OpenStudy (anonymous):

I Like you.You know yourself.Phillip charged $400 worth of goods on his credit card. On his first bill, he was not charged any interest, and he made a payment of $20. He then charged another $18 worth of goods. On his second bill a month later, he was charged 2% interest on his entire unpaid balance. How much interest was Phillip charged on his second bill?

OpenStudy (anonymous):

$7.96

OpenStudy (anonymous):

x = (398) (0.02) x= 7.96

OpenStudy (anonymous):

thanks,How did you solve it?

OpenStudy (anonymous):

Percentage = base x rate You're looking for the percentage. You're base is 398 - He spent 400, but paid 20. So he's left with 380. But he spent another 18 so a total of 398 unpaid. The rate is the %. Since it's 2%, divide 2 by 100, you get .02 Plug everything in... percentage = 398 * 0.02 percentage = 7.96 Hope I was clear enough. :)